# Covariances for Free: Exploiting Mean Distributions for Federated Learning with Pre-Trained Models

### v1

@@ 4.1 Motivation
Figure 2: Analysis showing increasing communication cost for Fed3R with increasing number of clients assuming 100 classes per client. This is due to the high dimensionality of the features ($d=512$ for SqueezeNet, $d=768$ for ViT-B/16 and $d=1280$ for MobileNetV2).
@@
Communication cost. While Fed3R is more effective than FedNCM, it requires each client to send $C$ vectors of size $d$ and a $d\times d$ matrix, significantly increasing the communication overhead by $d^{2}K$ compared to FedNCM which only shares the class means (see Table 1). Scaling the method to a large number of clients requires much higher communication cost, as it scales linearly with number of clients and quadratically with the feature dimension, as shown in Fig. 2. Smaller neural network models often have a very high-dimensional feature space. For instance, ResNet-50 has $d=2048$ with 25.6 million parameters, MobileNetV2 has $d=1280$ with 3.4 million parameters while ViT-B/16 has more parameters (86 million) with $d=768$. Considering cross-device FL settings (Kairouz et al., 2021), having millions of client devices, the communication cost needed for Fed3R would be enormous. In settings with low-bandwidth communication, using Fed3R is not realistic. See Appendix E for more discussion.
@@
Potential privacy concerns. Sharing only class means provides a higher level of data privacy compared to sharing raw data, as prototypes represent the mean of feature representations. It is not easy to reconstruct exact images from prototypes with feature inversion attacks, as shown by (Luo et al., 2021). As a result, sharing class means is common in many recent works (Tan et al., 2022b; a; Shysheya et al., 2022; Legate et al., 2023a). However, sharing feature matrices (Fanì et al., 2024) exposes the feature distribution of the client to the server since all clients employ the same frozen pre-trained model to extract features. Sharing covariances makes the clients more vulnerable to attacks if additional secure aggregation protocols are not implemented (Bonawitz et al., 2016).
While exploiting second-order statistics (using Fed3R (Fanì et al., 2024), for example) yields significant gains in accuracy as shown in Fig. 1, it faces the above mentioned issues. We propose instead to estimate class covariances at the server using only class means and counts from clients. This will allow us to exploit second-order statistics without actually sharing them between clients and server. Following (Legate et al., 2023a; Luo et al., 2021), we use class frequencies from clients [^1] Our method requires communication of class frequencies which could raise minor privacy concerns; in Appendix L we perform an extensive evaluation of potential methods to address those concerns. since it only quantifies the client data while not revealing any information at the data or feature level.
4.2 Estimating Class Covariances Using Only Client Means
We aim to use the pre-trained feature extractor and initialize the global classifier by exploiting second-order statistics for each class. We propose estimating the client class covariances using only class means. Our method leverages the statistical properties of sample means to derive an unbiased estimator of the class population covariance based only on class means. This estimator is then used to initialize the classifier (see Fig. 3 for an overview).
Figure 3: Federated Learning with COvariances for Free (FedCOF). Each client $k$ communicates only its class means $\hat{\mu}_{k,c}$ and counts $n_{k,c}$. On the server side, (A) we use a provably unbiased estimator $\hat{\Sigma}_{c}$ (denoted by solid lines) of population covariance $\Sigma_{c}$ (denoted by dashed lines) based on the received class means (see Section 4.2). (B) We initialize the linear classifier using the estimated second-order statistics and remove the between-class scatter matrix as discussed in Section 4.3.
@@ 4.2 Estimating Class Covariances Using Only Client Means
Assume that the features of a class $c$ are drawn from a population with mean $\mu_{c}$ and covariance $\Sigma_{c}$. The features computed by each client are a random sample drawn from this population distribution. Using the statistical properties of the sample mean we can prove the following proposition.
Proposition 1.
$\overline{F}_{k,c}=\frac{1}{n_{k,c}}\sum_{j=1}^{n_{k,c}}F_{k,c}^{j},$ (6)
is distributed with mean $\mathbb{E}[\overline{F}_{k,c}]=\mu_{c}$ and covariance $\mathrm{Var}[\overline{F}_{k,c}]=\frac{\Sigma_{c}}{n_{k,c}}$.
@@
In Appendix B we provide the proof of this well-known result about the distribution of sample means and covariances. Intuitively, since $\Sigma_{c}=n_{k,c}\mathrm{Var}[\overline{F}_{k,c}]$, this proposition suggests that by assigning multiple sets of $n_{k,c}$ features to a single client, we can compute the empirical covariance of the client’s class means over multiple assignments, providing an estimator of population covariance $\Sigma_{c}$.
@@
However, in federated learning data are assigned only once to each client, and there are $K$ clients in the federation, each with $n_{k,c}$ features and $n_{i,c}\neq n_{j,c}$ for $i\neq j$. To estimate the population covariance $\Sigma_{c}$, we need an estimator that accounts for the contributions of all $K$ clients. In the following proposition, we propose such an estimator.
@@
To prove that $\hat{\Sigma}_{c}$ is an unbiased estimator of the population covariance, we show in Appendix C that $\mathbb{E}[\hat{\Sigma}_{c}]=\Sigma_{c}$. Under the iid assumption of client feature distribution before federated training with a pre-trained model, the class features of each client can be considered as a random sample of size $n_{k,c}$, and the global class features as a sample of size $N_{c}$. By applying Proposition 1, we find that each client class mean has $\mathbb{E}[\overline{F}_{k,c}]=\mu_{c}$ and $\mathrm{Var}[\overline{F}_{k,c}]=\frac{\Sigma_{c}}{n_{k,c}}$, while the global class mean $\hat{\mu}_{c}$ has $\mathbb{E}[\hat{\mu}_{c}]=\mu_{c}$ and $\mathrm{Var}[\hat{\mu}_{c}]=\frac{\Sigma_{c}}{N_{c}}$. Using the properties of expectation, we complete the proof.
Covariance shrinkage. Van Ness (1980) and Friedman (1989) proposed adding an identity matrix to the covariance matrix to stabilize the smaller eigenvalues. Shrinkage helps especially when the number of samples is fewer than the number of feature dimensions resulting in a low-rank covariance matrix. Shrinkage has been adopted in recent works on continual learning using feature covariances (Goswami et al., 2023; Magistri et al., 2024). In our problem, the proposed covariance estimation using a limited number of clients may poorly estimate the population covariance $\Sigma_{c}$. So, we perform shrinkage to better estimate the class covariances from the client means as follows:
$\hat{\Sigma}_{c}=\frac{1}{K-1}\sum_{k=1}^{K}n_{k,c}(\hat{\mu}_{k,c}-\hat{\mu}_{c})(\hat{\mu}_{k,c}-\hat{\mu}_{c})^{\top}+\gamma I_{d}$ (8)
@@
Impact of the number of clients. The quality of estimated covariances depends on number of clients. More clients will give more class means, thus improving the estimate compared to fewer clients. While realistic FL settings has thousands of clients (Hsu et al., 2020; Kairouz et al., 2021), there can be FL settings with fewer clients. In cases with few clients, we propose to sample multiple means from each client to increase total number of means used for covariance estimation. This can be easily done by randomly sampling multiple subsets of features in each client without replacement and computing a mean from each of these subsets. We validate this in experiments (see Fig. 11(a)).
@@
The iid assumption. In FL each client has its own data, typically distributed in a statistically heterogeneous or class-imbalanced manner according to a Dirichlet distribution (Hsu et al., 2019). As a result, each client has data belonging to a different set of classes in varying quantities, resulting in non-iid data distributions across clients. However, note that the samples belonging to the same class in different clients are sampled from the same distribution. We exploit this fact in FedCOF. We later show empirically that our method can be successfully applied to non-iid FL scenarios involving thousands of heterogeneous clients on iNaturalist-Users-120K (Hsu et al., 2020). We hypothesize that this can be attributed to the strong generalization capabilities of pre-trained models, which help in moving the distribution of class samples across clients towards an iid feature distribution even if the class distribution across clients is non-iid (Nguyen et al., 2023). We analyze the bias of the estimator under non-iid assumptions in Appendix J and evaluate the performance of FedCOF in feature shift settings (Li et al., 2021) in Appendix K.
Algorithm 1 Federated Learning with COvariances for Free (FedCOF)
@@
  Input:
  $C$: set of all classes
  $f_{\theta}$: pre-trained feature extractor
  $n_{k,c}$: number of samples for class $c$ in client $k$
  $X_{k,c}$: samples of class $c$ in client $k$
@@
  Input:
  $\hat{\mu}_{k,c},n_{k,c}$ sent from $K$ clients
  $\lambda>0$, $\gamma>0$: hyper-parameters
@@
     $\hat{\mu}_{c}=\frac{1}{N_{c}}\sum_{k=1}^{K}n_{k,c}\hat{\mu}_{k,c};$ $N_{c}=\sum_{k=1}^{K}n_{k,c}$ class mean
[…]
@@
  $\hat{\mu}_{g}=\frac{1}{N}\sum_{c=1}^{C}N_{c}\hat{\mu}_{c}$       $N=\sum_{c=1}^{C}N_{c}$      # global mean
  $B=\left[\hat{\mu}_{c}N_{c}\right]_{c=1}^{C},$ Eq.(9)
  $\hat{G}=\sum_{c=1}^{C}(N_{c}-1)\hat{\Sigma}_{c}+N\hat{\mu}_{g}\hat{\mu}_{g}^{\top}$
  $W^{*}=(\hat{G}+\lambda I_{d})^{-1}B,$ Eq. (11)
@@
4.3 Classifier Initialization with Estimated Covariances
Having derived how to compute class covariances from client means, we now discuss how to use class covariances to set the classifier weights and then replace the empirical class covariances with our estimated class covariances, as illustrated in Fig. 3.
Proposition 3.
$B=\left[\hat{\mu}_{c}N_{c}\right]_{c=1}^{C},$ (9)
We prove this result in Appendix D. The proof is based on the key observation that $G=FF^{\top}$ from ridge regression is an uncentered and unnormalized empirical global covariance. By employing the empirical global covariance definition and decomposing it into within-class and between-class scatter matrices, we obtain the above formulation of $G$.
To analyze the impact of the two scatter matrices, we consider the centralized setting in Fig. 4 and empirically find that using only within-class scatter matrix performs slightly better than using total scatter matrix in Eq. 10. As a result, we propose to remove the between-class scatter and initialize the linear classifier at the end of the pre-trained network using the within-class covariances $\hat{\Sigma}_{c}$ which are estimated from client means using Eq. 8, as follows:
$W^{*}$ $=(\hat{G}+\lambda I_{d})^{-1}B;\quad\quad\hat{G}$ $=\sum_{c=1}^{C}(N_{c}-1)\hat{\Sigma}_{c}+N\hat{\mu}_{g}\hat{\mu}_{g}^{\top}.$ (11)
Theoretically, we observe that a similar approach is used in Linear Discriminant Analysis (Ghojogh & Crowley, 2019), which employs only within-class covariances for finding optimal weights.
Figure 4: Analysis showing improved accuracy by removing between-class scatter for classifier weights using Eq. 11 in centralized setting.
To summarize, we estimate the covariance matrix for each class using only the client means (Eq. 8) and use the estimated covariances to initialize the classifier as in Eq. 11. Finally, we normalize the weights for every class to account for class imbalance in the entire dataset. We provide the summary in Algorithm 1.
FedCOF with multiple rounds. While the proposed estimator requires class means from all clients in a single round, this might not be realistic in settings in which clients appear in successive rounds based on availability. In the case of multi-round classifier initialization (see FedCOF in Fig. 7 before fine-tuning), the server uses all class means and counts received from all clients seen up to the current round and stores the accumulated means and counts for future use. As a result, FedCOF uses statistics from all clients seen up to the current round, similar to Fed3R. Thus FedCOF converges when all clients are seen at least once.

### v3

@@ 4.1 Motivation
Communication cost. While Fed3R is more effective than FedNCM, it requires each client to send an additional $d\times d$ matrix, significantly increasing the communication overhead by $d^{2}K$ compared to FedNCM which only shares the class means (see Section˜1). Fed3R scales linearly with number of clients and quadratically with the feature dimension. FedPFT beitollahifoundation shares multiple sets of $(\mu_{k,c},\Sigma_{k,c})$ from clients, further increasing communication costs. Considering cross-device FL settings (kairouz2021advances, ), having millions of clients, the communication cost required for these methods would be enormous.
Recent works koo2024towards ; weng2024probabilistic focus on parameter-efficient federated fine-tuning where the per-round communication costs are significantly reduced, enabling applications in low-bandwidth communication settings. For example, recent work (koo2024towards, ) shows that using pre-trained language models (RoBERTa-base) can yield performance similar to full fine-tuning while using rank-1 LoRA updates and thereby reducing communication cost by 99.8%. Sharing a similar goal of reducing communication costs, we propose an unbiased covariance estimator without sharing client covariances for training-free FL.
@@
Potential privacy concerns. Sharing only class means provides a higher level of data privacy compared to sharing raw data, as prototypes represent the mean of feature representations. It is not easy to reconstruct exact images from prototypes with feature inversion attacks (luo2021no, ). As a result, sharing class means is common in many recent works (tan2022federated, ; tan2022fedproto, ; shysheya2022fit, ; legate2023guiding, ). On the other hand, Fed3R shows that sharing second-order statistics improves performance compared to sharing class means. However, this could expose the feature distribution of clients to the server since all clients employ the same frozen pre-trained model to extract features fani2024accelerating . Sharing covariances makes clients more vulnerable to attacks if secure aggregation protocols are not implemented (bonawitz2016practical, ).
4.2 Estimating Covariances Using Only Client Means
@@
Our method leverages the statistical properties of sample means to derive an unbiased estimator of the class population covariance based only on per-client class means (see Figure˜2). We model the global features of each class $c$ as drawn from a multivariate distribution with mean $\mu_{c}$ and covariance $\Sigma_{c}$, and assume that the local features for class $c$ computed by each client using the shared frozen pre-trained model are iid. Under this assumption, class features in a client form a random sample from the class population.
For client $k$, let $\{F^{j}_{k,c}\}_{j=1}^{n_{k,c}}$ denote the feature vectors of class $c$, where $n_{k,c}$ is the number of samples from class $c$ assigned to client $k$. The sample mean of these features, $\overline{F}_{k,c}=\frac{1}{n_{k,c}}\sum_{j=1}^{n_{k,c}}F_{k,c}^{j}$, is itself a random variable with expectation and variance given by:
$\mathbb{E}[\overline{F}_{k,c}]=\mu_{c},\quad\mathrm{Var}[\overline{F}_{k,c}]=\frac{\Sigma_{c}}{n_{k,c}}$ (4)
@@
This well-known result – whose proof we include in Appendix˜B for reference – implies that the variance of sample means over subsets of size $n_{k,c}$ reflects the underlying class covariance. In principle, by assigning multiple subsets of $n_{k,c}$ features to a single client and computing the empirical covariance of their means, one could recover $\Sigma_{c}$, since $\Sigma_{c}=n_{k,c}\mathrm{Var}[\overline{F}_{k,c}]$.
@@
However, in federated learning data are assigned only once to each client, and there are $K$ clients in the federation, each with $n_{k,c}$ features and $n_{i,c}\neq n_{j,c}$ for $i\neq j$. To estimate the population covariance $\Sigma_{c}$, we need an estimator that accounts for the contributions of all $K$ clients. In the following proposition, we propose just such an estimator.
@@
Proof. To prove that $\hat{\Sigma}_{c}$ is an unbiased estimator of the population covariance, we show that $\mathbb{E}[\hat{\Sigma}_{c}]=\Sigma_{c}$. Under the iid assumption of client feature distribution with a frozen pre-trained model, the class features of each client can be considered as a random sample of size $n_{k,c}$, and the global class features as a sample of size $N_{c}$. By exploiting the result in Equation˜4, each client class mean has $\mathbb{E}[\overline{F}_{k,c}]=\mu_{c}$ and $\mathrm{Var}[\overline{F}_{k,c}]=\frac{\Sigma_{c}}{n_{k,c}}$, while the global class mean $\hat{\mu}_{c}$ has $\mathbb{E}[\hat{\mu}_{c}]=\mu_{c}$ and $\mathrm{Var}[\hat{\mu}_{c}]=\frac{\Sigma_{c}}{N_{c}}$. Using this fact and applying the properties of expectation to $\mathbb{E}[\hat{\Sigma}_{c}]$, we complete the proof. In Appendix˜C we provide the detailed proof.
Figure 2: Federated Learning with COvariances for Free (FedCOF). Each client $k$ communicates only its class means $\hat{\mu}_{k,c}$ and counts $n_{k,c}$. On the server side, (A) we use a provably unbiased estimator $\hat{\Sigma}_{c}$ (denoted by solid lines) of population covariance $\Sigma_{c}$ (denoted by dashed lines) based on the received class means (see Section˜4.2). (B) We initialize the linear classifier using the estimated second-order statistics and remove the between-class scatter matrix as discussed in Section˜4.3.
Covariance shrinkage. ShrinkagenessOn ; friedman1989regularized stabilizes covariance estimation by adding a scaled identity matrix to the covariance matrices, which regularizes small eigenvalues and reduces estimator variance. This is particularly helpful when the number of samples is smaller than the feature dimensionality, leading to low-rank covariances, and has been recently adopted in continual learning methods leveraging feature covariances (goswami2024calibrating, ; magistri2024elastic, ). In the FL setup, the covariance estimation using a limited number of clients may poorly estimate the population covariance $\Sigma_{c}$. So, we perform shrinkage to better estimate the class covariances from the client means as follows:
$\hat{\Sigma}_{c}=\frac{1}{K-1}\sum_{k=1}^{K}n_{k,c}(\hat{\mu}_{k,c}-\hat{\mu}_{c})(\hat{\mu}_{k,c}-\hat{\mu}_{c})^{\top}+\gamma I_{d},$ (6)
@@
Impact of number of clients. The quality of estimated covariances depends on the number of clients. More clients will give more means and improve the estimate compared to fewer clients. While realistic settings have thousands of clients (hsu2020federated, ; kairouz2021advances, ), there can be FL settings with fewer. In such cases, we propose to sample multiple means from each client to increase number of means used for covariance estimation. This can be done by randomly sampling subsets of features in each client and computing a mean from each subset. We validate this approach experimentally (see Figure˜5). We discuss more on sampling multiple class means in Appendix˜E.
4.3 Classifier Initialization with Estimated Covariances
Having derived how to compute class covariances from client means, we now describe how FedCOF leverages the estimated class covariances to initialize classifier weights.
Proposition 2.
$B=\left[\hat{\mu}_{c}N_{c}\right]_{c=1}^{C},\quad G=\sum_{c=1}^{C}(N_{c}-1)\hat{S}_{c}+\sum_{c=1}^{C}N_{c}(\hat{\mu}_{c}-\hat{\mu}_{g})(\hat{\mu}_{c}-\hat{\mu}_{g})^{\top}+N\hat{\mu}_{g}\hat{\mu}_{g}^{\top},$ (7)
where the first two terms $G_{with}=\sum_{c=1}^{C}(N_{c}-1)\hat{S}_{c}$  and  $G_{btw}=\sum_{c=1}^{C}N_{c}(\hat{\mu}_{c}-\hat{\mu}_{g})(\hat{\mu}_{c}-\hat{\mu}_{g})^{\top}$ represents the within-class and between class scatter respectively, while $\hat{\mu}_{c}$, $\hat{S}_{c}$ and $N_{c}$, denote the empirical mean, covariance and sample size for class $c$, respectively.
Proof. The proof is based on the observation that $G=FF^{\top}$ from ridge regression is an uncentered and unnormalized empirical global covariance. By using the empirical global covariance definition and decomposing it into within-class and between-class scatter, we obtain the above formulation of $G$. In Appendix˜D, we provide the detailed proof.
Table 2: Analysis showing improved accuracy by removing between-class scatter for classifier initialization in centralized setting using pre-trained SqueezeNet.
Dataset $G_{btw}$ $G_{with}$ Acc.($\uparrow$)
CIFAR-100 57.1
57.3
ImageNet-R 37.6
38.6
CUB200 50.4
53.7
Stanford Cars 41.4
44.8
To analyze the impact of the two scatter matrices, we first consider the centralized setting in Table˜2 and empirically find that using only within-class scatter matrix performs better than using total scatter matrix in Equation˜7. We then analyze their spectral properties via the condition number, defined for a matrix $G$ as $\mathcal{K}(G)=\lambda_{\text{max}}(G)/\lambda^{+}_{\text{min}}(G)$, where $\lambda_{\text{max}}(G)$ and $\lambda_{\text{min}}^{+}(G)$ are the largest and smallest non-zero eigenvalue, respectively. For a SqueezeNet backbone in the centralized setting, we observe that $G_{btw}$ is highly ill-conditioned, with condition numbers $\mathcal{K}(G_{btw})$ of $3.0\times 10^{7}$, $5\times 10^{7}, 2.2$$2\times 10^{7}$, $1.3\times 10^{7}$ on CUB, Cars, ImageNet-R and CIFAR-100, respectively, while $G_{with}$ is much better conditioned ($\mathcal{K}(G_{with})$: $4.5\times 10^{3}$, $2.5\times 10^{4}$, $8.2\times 10^{2}$, $6.3\times 10^{3}$). Including $G_{btw}$ can cause numerical instability due to its poor conditioning, leading the classifier to overfit directions with small eigenvalues that capture noise or dataset-specific artifacts. This is consistent with the results in Appendix˜F, where we show that using $G_{btw}$ leads to higher overfitting on the training data.
As a result, we propose to remove the between-class scatter and initialize the linear classifier at the end of the pre-trained network using the within-class covariances $\hat{\Sigma}_{c}$ which are estimated from client means using Equation˜6, as follows:
$W^{*}$ $=\hat{G}^{-1}B;\kern 5.0pt\quad\hat{G}$ $=\sum_{c=1}^{C}(N_{c}-1)\hat{\Sigma}_{c}+N\hat{\mu}_{g}\hat{\mu}_{g}^{\top}.$ (8)
Theoretically, we observe that a similar approach is used in Linear Discriminant Analysis (ghojogh2019linear, ), which employs only within-class covariances for finding optimal weights. By removing between-class scatter, we propose a more effective classifier initialization than Fed3R (which uses $G$ from Equation˜7 and considers both within- and between-class scatter matrices).
We summarize in Algorithm˜1 how we estimate the covariance matrix for each class using only the client means and use the estimated covariances to initialize the classifier as in Equation˜8. The normalization of the weights accounts for class imbalance in the entire dataset.
@@ 4.3 Classifier Initialization with Estimated Covariances
Impact of the iid assumption. FedCOF initializes the classifier with the class covariance estimator (Equation 6), unbiased only under the assumption of iid pre-trained features per class (Section 4.2). In FL each client has its own data, typically distributed in a statistically heterogeneous or class-imbalanced manner according to a Dirichlet distribution (hsu2019measuring, ). As a result, each client has data from different set of classes, resulting in non-iid data distributions across clients. However, note that the samples belonging to the same class in different clients are sampled from the same distribution. We exploit this fact in FedCOF. We later show empirically that our method can be successfully applied to non-iid FL scenarios involving thousands of heterogeneous clients on iNaturalist-Users-120K (hsu2020federated, ). We hypothesize that this can be attributed to the strong generalization capabilities of pre-trained models. We analyze the bias of the estimator under non-iid assumptions for the same class and also evaluate FedCOF in feature shift settings (lifedbn, ) in Appendix˜G.
FedCOF in multiple rounds. While the proposed estimator requires class means from all clients in a single round, this might not be realistic in settings in which clients appear in successive rounds based on availability. For multi-round classifier initialization (see FedCOF in Figure˜3 before fine-tuning), the server uses all class means and counts received from all clients seen up to the current round and stores the accumulated means and counts for future rounds. As a result, FedCOF uses statistics from all clients seen up to the current round, similar to Fed3R. This can be easily implemented by ensuring on the client side that each client transfers statistics to the server only once, avoiding repeated transfer of statistics. FedCOF converges when all clients are seen at least once, with total communication cost equal to the single round-case (see Appendix˜H for details on communication cost).
Privacy aspects. FedCOF requires each client to send its class means and frequencies (see Eq. 5), similar to CCVR (luo2021no, ). While this avoids sharing feature-level data, it does not guarantee strong privacy and may introduce concerns, particularly because it complicates the secure aggregation protocol applicable in Fed3R. Nonetheless, as discussed in Appendix˜I, FedCOF can be combined with additional privacy-preserving mechanisms, such as adding noise to the shared statistics or adapting the method to support secure aggregation. The latter preserves FedCOF’s performance but increases communication overhead to the level of Fed3R.
Algorithm 1 FedCOF: Federated Learning with COvariances for Free
@@
  Input: $C$: set of all classes, $f_{\theta}$: pre-trained model, $X_{k,c}$: samples of class $c$ in client $k$, $n_{k,c}$: number of samples in $X_{k,c}$
@@
  Input: $\hat{\mu}_{k,c},n_{k,c}$ sent from $K$ clients, $\gamma>0$
@@
   $\hat{\mu}_{c}=\frac{1}{N_{c}}\sum_{k=1}^{K}n_{k,c}\hat{\mu}_{k,c};$ $N_{c}=\sum_{k=1}^{K}n_{k,c}$
   $\hat{\Sigma}_{c}\!=\!\frac{1}{K-1}\underset{k=1}{\overset{K}{\sum}}n_{k,c}(\hat{\mu}_{k,c}-\hat{\mu}_{c})(\hat{\mu}_{k,c}-\hat{\mu}_{c})^{\top}\!+\gamma I_{d},$ Eq.(6)
@@
  $\hat{\mu}_{g}=\frac{1}{N}\sum_{c=1}^{C}N_{c}\hat{\mu}_{c}$,         $N=\sum_{c=1}^{C}N_{c}$
  $B=\left[\hat{\mu}_{c}N_{c}\right]_{c=1}^{C},$    $\hat{G}=\sum_{c=1}^{C}(N_{c}-1)\hat{\Sigma}_{c}+N\hat{\mu}_{g}\hat{\mu}_{g}^{\top}$
  $W^{*}=\hat{G}^{-1}B,$ Eq. (8)